OpenStudy (anonymous):

2x/3-5x/9-x/6-5x/12=13/2

OpenStudy (anonymous):

Find the LCM of (3,9,6,12) 3/_3,6,9,12 2/_1,2,3,4 2/_1,1,3,2 3/_1,1,3,1 /_1,1,1,1 LCM is 3*2*2*3=36 x(24-20-6-15)/36=13/2 -17x/36=13/2 x=-36*13/17*2 x=-234/17